A 2030 kg space station orbits Earth at an altitude of 4.69×105 m. Find the magnitude of the force with which the space station attracts Earth. The mass and mean radius of Earth are 5.98×1024 kg and 6.37×106 m, respectively.
Answer #1

Force between earth and satellite is given by

F = G*m1*m2/r2

m1 = 2030

m2=5.98*1024

G= universal gravitational constant = 6.67*10-11

how to know r is tricky

altitude = distance form the surface of earth

r= distance between the Center of earth and satellite

so

r= altitude +radius of earth in meter

so r=4.69*10^5 +6.37*10^6

so after plug-in the values in formula we got

F =17311.62 N answer

or F = 1.73*104 N